- We are ready to implement ASS for PRM
I have added an IPC sender from the LSC model, to send POPDC to ASS. I have copied over the structure of the arms' ASS, to do the same for PRCL. I have set it up to dither the PRM, and feed back to the PRM. I did not include an LSC set, since I'm assuming that we'll set the input pointing with the arms, and just want to move the PRM to maximize POPDC.
Models have been compiled, installed, and restarted, and the daqd was restarted.
I'm not really sure why the ASS was involved in this. I feel like it might have been simpler to just do everything in the ASX model, to keep things cleaner. Also, the IPC blocks for this stuff (in both ASS and ASX) are not on the top level of the model. I had thought that this was expressly forbidden (although I'm not sure why). I'm emailing Jamie, to see if he remembers what, if anything, is breakable if the IPC blocks are down a level.
I'm not sure if it's forbidden by the RCG, but you should definitely NOT do it. All IO, whether it be between ADC/DACs or IPCs, should always be at the model top level. That's what keeps things portable, and makes it easier to keep track of where are signals are going/coming from.
Getting rid of the LO transmission will certainly help / be good. After adding these channels, the RFM model is regularly hitting 62usec (out of a max acceptable of 60).
In the past, we used to use Stefan's 'ezcademod' or Matt's 'ezlockin' to do auto phase adjustment.
JoeB / Jamie are working on python replacements for these tools, but in the near term possibly I can make a bash script to use ezcaservo and the existing LOCKINs to do this.
Over the last three days, I've had the interferometer to test and optimize the ASX Servo. Based on what I have seen, I think the conclusion is that with the current parameters, the servo does its job provided the input pointing set up at the endtable with the coarse adjustment knobs is reasonably good. Once the cavity is aligned and IR transmission maximized using ASS, I have been able to get the green transmission up to 0.8 which is close to the best we had pre-vent. I have not been elogging regularly over the last few days, so this one is going to be a longish one.
Major changes made:
Details of tests runs:
For the most part, I have been trying to optimize the PZT mirror dither servo. To this end, I did the following:
Attempt to measure transfer function:
One of the things that came up during my presentation was how fast the loop was capable of responding. I was able to get a quantitative idea of this by playing around with the overall servo gain. Initially, it took ~30 seconds for the servo to get the transmission up to its peak value, with a servo gain of 1. When I ramped this up to 5, the response was much faster, with the peak transmission being reached in ~5seconds.
I wanted to get a more quantitative picture, and hence tried to measure the transfer function by first injecting an excitation into the 'SIG' filter-bank in the demodulation stage. However, coherence between the IN1 and IN2 signals was very poor for all the amplitude configurations I tried. At Jenne's suggestion, I tried injecting the excitation at the control-filters stage, but found no improvement. Perhaps the amplitude envelope was wrong or the measurement technique has to be rethought.
New MEDM screen:
PRMI(sb) lock was recovered
- Stared at the time series data of the REFL demod signals, and decided to use REFL165I&Q for the locking.
- Jiggled the demodulation phase of REFL165 and POP110. Changed the servo gains.
- Finally found a short lock. Further optimized the parameters.
- PRM ASC was turned on by giving the identity matrices for the input and output matrices.
Now just hitting the up button is sufficient to engage the ASC servo.
- Under the presence of the ASC, the PRMI is indefinitely locked as before.
- Reacquisition is also instantaneous. (It acquires even if the ASC is left "on".)
- Actually the lock is somewhat robust even when the PRM ASC is not used.
This is VERY GOOD as we can skip one of the steps necessary for the full lock.
Although, the seismic on Friday night is very quiet.
The spot motion at POP seems to be somewhat pitch/yaw mixed, in stead of previous "totally-dominated-by-yaw" situation.
Demod phase adjustment
- Shook PRM at 580Hz / 100cnt
- Swept the demod phase of REFL165 such that the PRM peak is minimized in the Q signal
- Open DTT. Measured transfer functions between REFL165I and the Q signals of each PD.
- Minimized the PRCL signal coupling in the signals.
- The resolution of the adjustment was ~1deg.
Locking test with PRM/BS
Tried the lock acquisition only with PRM and BS. (cf. http://nodus.ligo.caltech.edu:8080/40m/8816)
This just worked nicely.
Today's locking parameters:
MC Trans: 17500
POP110I (in lock): 150
PRCL Source: REFL165(I) 106deg / 45dB / Normalization SQRT(10 POP110I) / Input MTRX 1.0
PRCL Trigger: POP110I x 1.0 50up 25down
PRCL Servo: G=+3.5 Acq: FM4/FM5 Opr: FM2/FM3/FM6/FM7
PRCL Actuator: PRM +1.0
MICH Source: REFL165(Q) 106deg / 45dB / Normalization SQRT(0.1 POP110I) / Input MTRX 1.0
MICH Trigger: POP110I x 1.0 50up 25down
MICH Servo: G=-10 Acq: FM4/FM5 Opr: FM2/FM3/FM6
MICH Actuator: (ITMX -1.0 / ITMY +1.0) or (BS 0.5 / PRM -0.267)
Now the SRM Yaw bias in yaw is functional without any strage behavior.
The problem was found at the connector of the flat ribbon cable from the DAC to the cross connect.
I used the extender board to diagnose the SRM coil driver circuit at 1X4.
The UL coil input did not show any sign of voltage no matter how the bias slider was jiggled.
I opened the side panel of the rack and found the signal was absent at the cross connect which relays two flat ribbon cables
for the SRM coil driver. I checked the DAC output with a multimeter. All the bias outputs were OK at the DAC.
Then I opened the IDC connector at the DAC side of the crossconnect as the signal was already missing there.
I found that the flat ribbon cable was a half line shifted from the supposed location.
This resulted a short circuit of the DAQ +/- pins for the SRM UL coil.
I recrimped the connector and now the SRM Yaw slider is back.
This changed the nominal position of the SRM. The new slider values were saved.
Notes to the fiber team:
I am aligning beam onto the RFPDs (I have finished all 4 REFL diodes, and AS55), in preparation for locking.
In doing so, I have noticed that the fiber lasers for the RFPD testing are always illuminating the photodiodes! This seems bad! Ack!
For now, I blocked the laser light coming from the fiber, did my alignment, then removed my blocks. The exception is REFL55, which I have left an aluminum beam dump, so that we can use REFL55 for PRM-ITMY locking, so I can align the POP diodes.
EDIT: I have also aligned POP QPD, and POP110/22. The fiber launcher for POP110 was not tight in its mount, so when I went to put a beam block in front of it and touched the mount, the whole thing spun a little bit. Now the fiber to POP110 is totally misaligned, and should be realigned.
What was done for the alignment:
1. Aligned the arms (ran ASS).
2. Aligned the beam to all the REFL and AS PDs.
3. Misaligned the ETMs and ITMX.
4. Locked PRM+ITMY using REFL11.
The following were modified to enable locking
(1) PRCL gain changed from +2.0 to -12.
(2) Power normalization matrix for PRCL changed from +10.0 to 0.
(3) FM3 in PRCL servo filter module was turned OFF.
5. POP PDs were aligned.
- IPANG aligned on the QPD. The beam seems to be partially clipped in the chamber.
- Oplev of the IFO mirrors are aligned.
- After the oplev alignment, ITMX Yaw oplev servo started to oscillate. Reduced the gain from -50 to -20.
I used MC_L signal from the Mode Cleaner as the desired signal with GUR2_X as witness signals. I observed good subtraction where coherence is high. But there was noise added in other frequency bands. I am not sure how to avoid that.
Please find attached documents that contains relevant plots.
An exercise of optimally subtracting one seismometer signal by another using weiner filters was done. Results have been summarized document attached.
I was receiving missing path error when I was trying to measure the MC spot positions. Jenne pointed out that Koji had moved all the unused scripts in scripts/ASS to /scripts/ASS/OBSOLETE yesterday and in the process one of the scripts that the MC spot position measurement script calls for (MeasureSpotPositions.py) must have also been moved to the OBSOLETE directory. I moved the script to /scripts/ASS/MC so that we know the script is being used and also changed its path in the main script.
1, Vacuum envelope grounds must be connected all times! After door removal reconnect both cables immediately.
2, The crane folding had a new issue of getting cut as picture shows.
3, Too much oplev light is scattered. This picture was taken just before we put on the heavy door.
4, We were unprepared to hold the smaller side chamber door 29" od of the IOC
5, Silicon bronze 1/2-13 nuts for chamber doors will be replaced. They are not smooth turning.
TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.
I have done some preliminary testing of the X-End Green ASS Servo. I will write a more detailed elog about this soon, but I thought I'd note down the important stuff here.
Yesterday, while we were venting, I aligned the X-arm to the green using the sliders on IFOalign, maximizing the transmission. Then I retook a power spectrum so as to determine the LO frequencies. Jenne pointed out that LO frequencies should not be integers (it usually suffices to append a .098725 or something to the frequency) so I made the necessary changes.
I did a first run of the servo yesterday, and more runs today. Notable points:
More details to follow.
While Gautam is working on the Xarm green ASS...
The EPICS monitor points for the ASS actuators were added to the ASS model.
This will be used for the offloading the ASS actuations to the alignment biases.
As this modification allowed us to monitor the actuation apart from the dithering,
now we can migrate the ASS actuation to the fast alignment offset on the suspension.
This modification to the offset moving scripts were also done.
We have looked a little more at the SRM situation. We aligned the SRM, and then aligned the oplev, so that we had a convenient monitor of the optic's motion.
When we use the _COMM channels, which are the usual ones on the IFO_ALIGN screen, the pitch slider makes pitch motion, but the yaw slider makes the oplev spot move ~45degrees from horizontal.
However, when we use the bias channels that are in the front end model, parallel to the ASC path, pitch moves pitch, and yaw moves pure yaw.
So, we conclude that the SRM coils are fine, and there is something funny going on with the slow part of the actuation.
Koji restarted the slow computer susaux, and burt restored it, but that did not fix the situation. We went inside and looked at all of the ribbon cable connections, and pushed them all in, but that also has not fixed things.
We have been looking at D010001-b, the coil driver board, and we think that's where the summing resistor network between the slow bias slider, and the coil outputs from the fast model exists. (It's not 100% clear, but we're confident that that's what is going on).
Tomorrow, we will pull the SRM's coil driver board, and see if any of the components in the slow slider path, before the summing point, look burned / broken / bad.
I have furthered Koji's work, and moved the filter on/off state for all the filter banks also to the burt snapshot.
Turning on the ASS is now much faster than it was originally, with the ezcawrites in series.
For the photodetector frequency response project, our new RF Switch Chassis (NI pxie-1071) arrived today. I took the switches out of the old chassis (Note for future generations: you have to yank pretty darn hard) and put them in the new chassis, which I mounted in rack 1Y1 as pictured.
The point of this new chassis is that its controller is compatible with our control room computer setup. We will be able to switch the chassis using TCP/IP or telnet, aiding in our automation of the measurement of photodetector frequency response.
Where are the foreline pressure gauges? These values are not on the vac.medm screen.
The new tip seal dry pump lowered the small turbo foreline pressure 10x
TP2fl after 2 day of pumping 65mTorr
I moved bunch of ezcawrite from the ASS Dither On script to a snapshot file.
This accelerated a half of the "up" time but still switching part is not in the snapshot.
If you find anything wrong with ASS, please notify me.
Cold cathode gauge CC1 -h (horizontal) just coming on 9.2e-5 Torr
P2 is the fore line pressure of the maglev. One can see the 4 Torr load during switching over to turbo pumping.
CC4 5e-9 Torr is the performance of the maglev pumping on the RGA only.
The annuloses are not pumped now. They are closed off to see how much outgassing plus leak they have.
Configuration: vacuum normal, annuloses not pumped
Precondition: 14 days at atm, IOO chamber north door was taken off as a new entrance, the ETMX chamber was not opened.
What is new in the vacuum system: new P1 pirani gauge, gold plated clean allen wrench and ..........what else was dropped?
Note: the wireless laptop did not fail once all day yesterday. I want to give credit to the person who is responsible for this.
I have a concern about the SRM suspension. The yaw alignment bias produces huge pitch coupling.
This could be a connector issue or the rubbing of the mirror on the EQ stops.
We have the photos of the magnets and they were not touching the OSEMs.
[Koji, Manasa, Jenne]
The Y arm was locked in IR, and we saw flashing in the Xarm (Gautam had the Xarm for green work when we began). I checked IPANG, and the beam was beautifully unclipped, almost perfectly centered on the first out of vacuum mirror. I aligned the beam onto the QPD.
We then swapped out the MC Y1 that we use at low power, and replace the usual 10% BS, so that we wouldn't crispy-fry MC REFL. Manasa adjusted the half wave plate after the laser, to maximize the power going toward the PMC. We relocked the PMC, and see transmission of ~0.84, which is at the high side of what we usually get. The beam was aligned onto MC REFL and centered on the WFS, and the MC was locked at nominal power. Koji tweaked up the alignment of the MC, and ran the WFS offset script. I aligned beam onto POP QPD and POP110 coarsely (using a flashing PRC, not a locked PRM-ITMY cavity, so the alignment should be rechecked). The arms have both been locked and aligned in IR....the green beams need to be steered to match the current cavity axis.
The AS beam, as well as REFL and POP, are all coming out of the vacuum nicely unclipped.
Notes: When Koji was aligning the SRM to get the SRC cavity roughly aligned (the AS flashes all overlapping), we noticed that there is some major pitch-yaw coupling. Serious enough that we should be concerned that perhaps some connector is loose, or an actuator isn't working properly. This should be checked.
Moral of the story: Coarse alignment of all mirrors is complete after pump-down and we have IR locked and aligned to both arms at nominal power.
Still to do:
* Restore PRM, align beam onto the REFL PDs.
* Lock PRM-ITMY cavity, align beam onto POP PDs.
* Align AS beam onto AS55.
* Recenter all oplevs.
* Recenter IPPOS and IPPANG at nominal power.
* Start locking!!
IFO P1=1mT PSL shutter is opened
I redid the power spectrum measurement for the X-arm green transmission after aligning the arm to green using the ITMX/ETMX Pitch and Yaw sliders on IFOalign.
The Y-axis now reflects the relative intensity noise (RIN), which I obtained by taking the average value of the X-arm green transmission using tdsavg. Based on this measurement, I have now picked four new frequencies at which to try and modulate the PZT mirrors: 10, 19, 34 and 39 Hz. Bandpass filters in the LIA stage have been appropriately modified.
For the photodetector frequency response project, I finished the construction of our baluns chassis and mounted it in rack 1Y1 (1st picture).
After consulting with Jenne, I mounted the fiber launcher for REFL165 on the AS table such that it would not cause an obstruction. I aligned the launcher using a multimeter to monitor the DC output of REFL165, but looking at the data I got, it seems I need to do a better alignment/focusing job to get rid of a bunch of noise.
The pumpdown has started at 8:38am
Manasa was here to confirm good alignment of the IFO
Inner jam nuts of AC bellow were torqued to 45 ft/lbs and door nuts were check on opened chambers.
Annulos were roughed down to 500 mTorr
Oplev servos turned off, PSL and green shutters closed before pumpdown started
I did an alignment check of the IFO before we start pumping down.
Arms were locked. PRM and SRM were aligned. Green was aligned to the arms for reference during the pump down.
Steve! It's a GO!
MC spot positions:
Oplevs were all centered yesterday and haven't drifted much. So I left them as is.
QPDs (IPPOS aligned from yesterday.
- I suppose the green transmission paths were thoroughly inspected and aligned in prior to the measurement
- Of course it is a BAD idea to use 60Hz as the LO frequency.
- Power spectra should be plotted as "RIN (relative intensity noise)" as the DC of 1 and 100 gives you 100 times different power spectra for the same beam.
Don't forget to subtract the offset from your DC values.
I had prepared two more PZT mounted mirrors for the Y-end some time back. These are:
I used the same QPD set-up and the methodology described here to do a full-range calibration of these PZTs. Plots attached. The calibration constants have been determined to be:
CVI-pitch: 0.316 mrad/V
CVI-yaw: 0.4018 mrad/V
Laseroptik pitch: 0.2447 mrad/V
Laseroptik yaw: 0.2822 mrad/V
CVI YAW CVI PITCH
Laseroptik YAW Laseroptik PITCH
In order to decide what frequencies to dither the 4 degrees of freedom (M1-pitch&yaw, M2-pitch&yaw) at, I took the power spectrum of the X and Y-arm green transmission (C1:ALS-TRX_OUT, C1:ALS-TRY_OUT). Plots showing the power spectra are attached. Looking at the power spectra, I would think that for the X-arm, it would be okay to dither at 40, 50, 60 and 70 Hz. In order to check if the piezos could respond to these frequencies, I used my QPD setup and shook the PZTs with a 100Hz, 1Vpp sinusoid, and saw that the spot moved smoothly on the QPD.
As for choosing the modulation amplitude, I did a simplistic approximation assuming that the misalignment only rotates the beam axis relative to the cavity axis, and determined what angle coupled 10% of the power into the next eigenmode. Assuming that this is small enough such that if we are already locked to TEM00, the dither won't kick it up to some higher-order mode, the LO amplitude should be in the range of 30-60 digital counts (determined using the PZT calibration constants determined here. This corresponds to a sine-wave of ~50mV amplitude reaching the PZTs (after HV amplification). I am not sure if this is too small, but according to the PZT datasheet, these platforms are supposed to have a resolution of 0.02 urad, which would correspond to the input signal changing by ~0.1 mV, so this signal should be capable of dithering the tip-tilt.
I have already added band-pass filters centered at these frequencies to the model (with a passband of 5Hz, 2Hz on either side), and low-pass filters to pull out the DC component of the output of the lock-in amplifiers. It remains to tune the gains of the filter stages. These parameters (frequency, amplitude of the LOs) may also have to be changed after tests). Hopefully the PZTs can be plugged in tomorrow, and I can try and make a measurement of the output matrix.
Koji also suggested that it may be good to have a path in the model that feeds back to the PZTs by dithering the cavity mirrors as opposed to the PZT mounted mirrors. I will work on incorporating this into the SIMULINK model (c1asx.mdl) and also into the master medm screen.
Power Spectra of Arm Green Transmission:
GET CAMERA IMAGES OF EVERYTHING
Check all OpLevs centered, in and out of vacuum
IPANG needed to be re-aligned today. Heavy doors are in place and bolts tight (torque 25 & 45).
Steve! We are ready for pump down!
I will check the IFO alignment once again early tomorrow morning before Steve starts pumping down.
I goofed on the transfer function requirement by not giving you the plant transfer function, which looks to be about 0.014 V/V, independent of frequency (PSL:1278). This needs to be compensated for in the electronic transfer function.
RGA background at day 12 of this vent . The maglev is pumping on the rga through VM2
[Manasa, Koji, Jenne]
We went into the BS and IOO chambers, and aligned the green beams such that they came out of the vacuum chamber. The idea here was to get the beams at the same height, but slightly offset in yaw. This required moving the Periscope on BS table, PBS in front of that periscope, the Periscope on the IOO table, and 2 steering mirrors on the IOO table after the 2nd periscope. The tables were not releveled, although we have aligned the full interferometer to this situation, so we do not want to touch the tables. The MC spot positions are still consistent with those measured earlier this afternoon, before this work, so I'm not concerned.
We confirmed that both green beams are hitting a good place (centered in pitch, and just left and right of center in yaw) on the mirror in the OMC chamber, and are getting to the center of the first mirror on the PSL table. We then coarsely aligned the beams on the PSL table.
We then relocked and aligned the arms for IR, and checked that the AS beam is centered on the mirrors in the BS chamber, and that the beam is coming out, and to the AS table. I touched the last mirror before PZT3 a small amount in yaw, and then PZT3 in pitch and yaw, until we saw the beam recentered on the first mirror on the AS table. At that point, we were also back to the center of the AS camera (which is good, since Koji had aligned all of that the other day). So, the AS beam is good.
We checked IPPOS, and have centered the beam on all the mirrors, and aligned the beam onto the QPD.
We checked IPANG, by looking through the viewports at the mirrors in the ETMY chamber. We are now centered in yaw, but clipping a bit low. This is what we want, since we always end up drifting high during the pump-down.
We see a nice, unclipped REFL beam on the camera.
We see a beam a little high on the POP camera, but Koji looked on the table with a card, and saw the beam....we just need to do minor alignment on the out of vac mirrors.
We checked again that the green TEM00 beams from both arms come to the PSL table.
We are getting POX and POY out, since we are using them to lock and align the arms.
Manasa and Koji recovered one clean allen key from the bottom of the chambers, but one remains, as a sacrifice to the vacuum gods.
I believe that, with the exception of checking the oplevs and taking photos of PR3, and the green steering optics, we have finished all of our vent tasks. We should do a quickie alignment on Monday, check the oplevs, take some photos, and put on the heavy doors. Pumping can start either Monday afternoon or Tuesday morning.
Attached is the finalized schematic. The general circuit topology should remain the same from this point forward, although individual component values are subject to change. I will also be adding some more annotations to ensure everything on the board is clear.
In general, I have finally included all of the correct components (i.e. front panel switches are now actually switches and front panel LEDs are now included). I also added an external 'Boost' switch, which can be used to enable or disable the boosts. The motivation for including this switch is that one might want to test functionality of the ISS without using the 'fancy' RMS detection and triggering circuitry. Additionally, one can disable the boosts when all the circuitry is stuffed in order to troubleshoot, so it essentially grants the board some flexibility in its operation.
I am now working on the PCB layout and I should hopefully have that done next week.
[Jenne, Manasa, Koji]
Earlier today, we locked and aligned both the X and Y arms.
I then went into the BS chamber, put on the BS' aperture, and put an aperture along the AS path. (We had Michelson fringes, so I centered the aperture around the fringes. I used one of the brass ruler things that we use to center the beam on ITMs and ETMs, on a riser. I put this aperture at the edge of the BS table, after the AS beam is launched toward the OMC chamber. The idea was to replace PR3 such that I could get the beam back through the BS aperture, and the brass ruler aperture, in hopes that we would see arm flashes, and not have to open the ITMY and ETMY heavy doors.)
I set references on the table so that I could put PR3 back in its original position, then removed PR3 from the chamber.
Steve set up a HeNe for me, that we pointed through the optic. The ghost beam was very high, indicating (as expected) that the wedge was not perfectly horizontal.
I took the suspension off of the cage and laid it down, as I have in the past.
I removed the optic from the suspension, to try to figure out which was the fat vs. skinny side. I noticed that there are very faint marks on the actual fat and skinny sides of the optic. (Mpral - for the LaserOptik mirrors, look for the faint lines that are the full width of the barrel, not the placement of the arrow which marks the HR side). I put the optic back in (HR side toward the back, fat side on the left (as you look at the face of the optic), which is consistent with the picture in the Optical Layout page of the Wiki, near the bottom.) the optic holder ring.
I put the suspension back on the cage, and saw that the HeNe's ghost beam was now nearly horizontal relative to the straight-through beam. Excellent. Also, the pitch balancing didn't seem to change noticably, which I determined was within "poking" distance of where we need it to be.
I put PR3 back onto the BS table, and adjusted it around until I got the beam through both the BS aperture, and the one on the AS path. As usual, this took quite a while, but as soon as I got through both of those apertures (really at the same place, not close to being through them, but as close as I could tell by eye - this is what took forever), Koji and Manasa saw flashes in the Yarm! Yay!
Since I had to move PR3 in angle a tiny bit, I reset the references, then dogged down PR3. We still had flashes, this time in both arms, so we closed up the light doors.
We have now locked and aligned both arms in IR after the adjustment of PR3, and see both arms' green at 01 or 02. We are about to start checking the green positioning on the periscopes. We will also need to check the AS path, as well as IPPOS and IPANG before we close up. We see REFL on the camera.
Separately - Manasa remembered that 2 clean things were dropped yesterday - a screw, and an allen key. Since they're both Clean, we're not too worried, although she thinks a long-armed person may be able to reach the allen key.
In PSL elog 1270, Evan elucidated the explicit requirements for the CTN ISS board. Essentially, the transfer function of the ISS should be something like:
TF_mag = (Unstabilized RIN) / (Calculated RIN Requirement)
I took Evan's data and did exactly this. I then designed a servo (using the general design I proposed here) to meet this requirement with a safety factor of ~10. By safety factor, I mean that if the ISS operates exactly according to theory, it should suppress the noise by a factor of 10 more than what is necessary/set out by the requirement. Below is a plot of the loop gain obtained directly from the requirement (the above expression for TF_mag) and the transfer function of the servo I am proposing.
I don't have the actual schematics attached as I was working with a LISO file and have yet to update the corresponding Altium schematic. The LISO file is attached and I will add the schematics later, although one can reference the second link to find a simple drawing.
[EricQ, Koji, Manasa]
We opened the BS chamber to check the status of the green beams. The X green has 3 steering mirrors before they hit periscope1 and the Y green transmits through all the optics giving no way to steer it.
We agreed to start fixing the Y green. The wedge angle of PR3 is steering the transmitted beam away in both pitch and yaw. Since we are restricted only to yaw movement (done by moving the periscope), we want he wedge angle to be oriented in the yaw as well.
Right now, the wedge is oriented at about 20-30 deg off (The mark on the side of the mirror does not indicate the wedge). So we see a pitch as well as yaw misalignment in the transmitted beam. The pitch misalignment is making the beam fall off the mirrors in periscope2.
We have decided to get the wedge angle set right for PR3 and redo the alignment for IR. Once we are aligned for the IR, we will modify the green layout.
The following slow channels have been added and are now being recorded by FB.
In order to integrate the data collected by the Raspberry-Pi from the Y-end doubling oven temperature controller and also the data from the frequency counter which will be hooked up to monitor the beat frequency, Koji helped me set up some slow EPICS record channels (in ALS as we felt this was most appropriate). The procedure for setting up slow channels was as follows (virtually identical to what is detailed in this elog:
I will now integrate these channels into my scripts, and make some simple MEDM screens.
I have made some minor changes to the model, made all the MEDM screens, and linked monitors on these to the appropriate channels. I have borrowed heavily from the C1ASS MEDM screens (particularly for the small filter modules-it was convenient to just copy and paste an existing module, and edit the channel names using EMACS/GEDIT), and have edited these to suit the needs of this servo. Some features:
I think I am now ready to take some measurements and try and optimize this servo. There is no green transmission at the PSL table at the moment, so not much can be done, though the first step would be to take the power spectrum of the error signal, which would help me decide the appropriate frequencies for the LOs. I would then have to add the appropriate filters to the model. The last, and most difficult step, would be the measurement of the output matrix, though Koji has given me some ideas about how this measurement can be done. I also have a template script ready, though I will only finalise this after optimising the servo and running it a couple of times manually.
Attached are screenshots of the MEDM screens.
We used our setup from yesterday (elog #8940) to measure transimpedance measurements for AS55, REFL11, REFL33, and REFL55, using our Newport 1611 FC-AC as reference. We connected the fibers to their respective telescopes such that the beams focused on their photodetectors, using a multimeter to maximize photodetector DC output. Plots are attached. At first glance, the poles seem to be where they're supposed to be.
Note that the procedure used today is similar to what the eventual automated procedure will be. The main differences are (1) The RF Switch will be used rather than manual switching (2) NWAG4395A will be used to collect data rather than netgpibdata (3) Data will be fit using vectfit4.m and compared to some canonical set.
Today I spent some time mounting the launcher and performing the same data collection for POX11. I think I still need to focus the launcher so the photodetector gets a good signal, but the data from today wasn't too bad. Additionally, I worked on matlab scripts to improve PDFR data analysis.
This time I collected data from the network analyzer using NWAG4395A in the netgpibdata directory. The advantage of this is that the computer tells the network analyzer to perform the sweep as well as retrieving the data.
For analysis, I improved my implementation of vectfit4.m so that it focuses in on the particular photodetector's predicted peaks and thus ignores much of the noise, giving a better fit. The raw data is the red circles in the 2nd attachment, while the fit is the blue line. I also had the program return the frequency value of the peak. For POX11, this was 1.106e+07 Hz.
I also finagled copies of existing programs to enable one to plot multiple transfer functions on the same axes. This function is /users/alex.cole/plottwo.m. I will eventually use this to compare new data to some canonical data so that we may monitor photodetector performance over time.
The eventual plan is to generate two plots per photodetector, one of which will compare new data to the canonical set, the other of which will show the fit of the data. Both will have subplots that zoom in around regions of interest (known peaks and notches), and the plot which displays the canonical set will also have Q's of peaks and their locations.
1)Power to the seismometers were turned down,
2)Guralp2 was moved to North side of POX table
3)Guralp2 was aligned in N-s Direction and leveled before connecting
4)Power to seismometers was turned on once Guralp2 was connected
We missed to check that we had the green transmitted to the PSL after flipping the SRC and PRC folding mirrors.
There is no green transmission reaching the PSL even after locking the arms to green.
We should fix this tomorrow. The BS heavy door should come off.
Steve! Do not start pump down tomorrow !
We closed the chambers last night with heavy doors and reopened it today.
I don't know what's going on here (why the channels are white), and I don't yet have a suggestion of where to look to fix it but...
Is there a reason that you're making a new model for this? You could just use and existing model at c1iscex, like the c1scx, and put your stuff in a top-names block. Then you wouldn't have to worry about all of the issues with adding and integrating a new model.
Koji just fixed this.
It seems that the new model's channels were not automatically added to the master file in the framebuilder (/opt/rtcds/caltech/c1/target/master). Adding the following two lines to the master file fixed the problem;
The box is now green. It looks like C1ASX.ini is created automatically in /opt/rtcds/caltech/c1/chans but the master file needs to be manually edited. The channels are now showing up on dataviewer etc. I have updated the information on the wiki page.
I have made a new model for the endtable PZT servo, and have put it in c1iscex. Model name is c1asx. Yesterday, Koji helped me start the model up. The model seems to be running fine now (there were some problems initially, I will post a more detailed elog about this in a bit) but some channels, which are computer generated, don't seem to exist (they show up as white blocks on the MEDM GDS_TP screen). I am attaching a screenshot of the said screen and the names of the channels. More detailed elog about what was done in making the model to follow.
C1:DAQ-DC0_C1ASX_STATUS (this is the channel name for the two leftmost white blocks)
These are roughly the steps I followed in setting up the new model for the endtable PZT servo - C1ASX.
I made a SIMULINK model of the servo, using MATLAB R2013a. The path to the model is /opt/rtcds/caltech/c1/userapps/release/isc/c1/models/c1asx.mdl. I am listing the parameters set on the CDS_PARAMETERS block:
Making, Compiling and Installing the Model:
After saving the model, I ssh-ed into c1iscex and ran the following commands:
rtcds make c1asx - this gave me a whole bunch of errors initially, which I tracked down to a naming problem in some of the from and goto flags: there should not be any spaces.
rtcds install c1asx
rtcds start c1asx - this gave me an error which said something like 'can't start/stop model.' Koji pointed out that given that a new model is being started, there is an additional step involved, which is to add the model name to the rtsystab file (this is located at /diskless/root/etc/rtsystab on framebuilder, and is mirrored in the various computers. It would be advisable to make sure that the changes are mirrored in the corresponding file on the computer in which the new model is being installed).
After adding the model name to the rtsystab file, I tried running rtcds start c1asx again. This time, no errors were output, but the model was not up and running as verified by looking at the C1:ASX_GDS_TP medm screen.
Koji suggested making a simple model (1 CDS parameters block, 1 ADC block and 2 filter modules, appropriately terminated) and see if that starts up, which it did. I then tried adding my servo minus the DAC block and recompiled and restarted the model. This too worked fine. I figured that the next logical step would be to add the DAC block to the model, and restart the model. But when I tried this, c1iscex crashed .
Jenne helped in restoring things to a working state (we reverted the c1asx model to just 2 filter modules, and went to the X-end and restarted the computer. This did not work the first time so I went back in and restarted it again, at which point we were able to ssh into c1iscex again and restart the four models running on it).
Since Manasa and Koji were working on getting things set up for the pumpdown,I did not try anything again till later in the evening, when Koji helped in debugging the problem further. In the meantime, at Jenne' suggestion, I made the model once again in MATLAB R2010b. In the evening, when I tried restarting the model, Koji suggested that the DAC channels in c1asx may be used by other models, at which point I realised I had set up excitation points on channels 8 through 15 of the DAC in c1scx (detailed here) in order to test the hardware at 1X9. I removed the excitation points from channels 8-11 of the DAC block in c1scx (these are the ones used in c1asx), and recompiled and restarted c1asx (using the above sequence of commands). I then tried recompiling and starting c1asx once more, and this time, it worked . At least, the GDS_TP screen suggests that the model is running alright, except for the fact that some computer generated channels seem to be missing. This problem is unresolved for now, and probably has something to do with the fact that C1ASX channels do not appear in Dataviewer.
I do not think this has to do with restarting framebuilder (I did the usual telnel fb 8088 followed by shutdown). In any case, I have added the new model to the CDS_FE_STATUS screen, and will continue to debug the same. I have also got a template medm screen (work in progress) which I will elog about soon as I get it done.
Note to self: There are 4 more excitation channels still hooked up to the DAC (channels 12-15) in the c1scx model. I plan to remove these and put them in c1asx.
We missed to check that we had the green transmitted to the PSL after flipping the SRC and PRC folding mirrors.
There is no green transmission reaching the PSL even after locking the arms to green.